# Cavity-less Brillouin strong coupling in a solid-state continuous system

### v1

@@ I Introduction
The particular case of light interacting with traveling acoustic waves, described by Brillouin-Mandelstam scattering, has received interest in recent years, with applications such as sensing [34, 35, 36], lasing [37, 38, 39] or integrated photonics [40, 41, 42]. A wide range of optoacoustic platforms is available, both in material [43, 44, 45] and structural [46, 47, 48, 49, 50, 51] diversity. Waveguides, with their defined direction of propagation, allow to have temporal control over the propagation of optical signals through them, which has in particular been exploited for signal processing [52, 53, 54, 55, 56, 57] and storage [58, 59, 60, 61]. The natural symmetry-breaking of Stokes and anti-Stokes processes has been exploited for active cooling [62, 63], without the need of working in the resolved sideband regime like in resonators [64, 65]. The shape of the spectrum of acoustic modes present in a waveguide is moreover crucially different from that of a resonator. In an optomechanical cavity, narrow discrete resonances are addressed, contrary to the close to continuous spectrum of modes present in a waveguide.
FIG. 1: (A) Sketch of an anti-Stokes Brillouin-Mandelstam scattering process in the strong coupling regime. (B) Dispersion diagram for the different modes involved in an anti-Stokes Brillouin-Mandelstam scattering process. Adapted from [77, 68]. (C) Schematic of a pump-probe experiment. (D) Schematic of a forced detuning experiment.
@@
FIG. 2: (A) Anti-Stokes Brillouin resonance as function of pump power. (B) Simulation of cavity-less optoacoustic strong coupling [66], showing good agreement with the experiment. (C) and (D) Optical anti-Stokes resonance in the strong coupling regime. (E) Optical anti-Stokes resonance in the weak coupling regime. Experimental data shown as a dark continuous line, theoretical data shown as a light dashed line.
@@
For low input powers, the system is weakly coupled ($g_{{\rm om}}$ = 13 ± 2 MHz) and the anti-Stokes resonance has the shape of a single Lorenzian centered at $\Omega_{\rm B}$ = 9.5997 ± 0.0001 GHz (Fig. 2 (E)). As the pump power increases ($g_{{\rm om}}$ = 57 ± 2 MHz), the system enters the strong coupling regime and the resonance splits (Fig. 2 (D)). Continuing to increase the pump power pushes the system deeper into the strong coupling regime ($g_{{\rm om}}$ = 70 ± 2 MHz, Fig. 2 (C)), resulting in a maximum measured split of $\delta_{{\rm split}}$ = 26.7 MHz for $g_{{\rm om}}$ = 71 ± 2 MHz. The shape of these two peaks is asymmetric, with the outer slopes of each peak following the shape of a Lorentzian slope, while the inner slopes are sharper. In order to validate our results, we compare with our theoretical model, which shows good agreement with the experiment (Fig. 2 (B)).
FIG. 3: (A) Diagram of the setup to perform a forced detuning measurement. CW: continuous wave; EDFA: Erbium-doped fiber amplifier; BPF: band-pass filter; E(A)OM: electro(acusto)optical modulator; ESA: electrical spectrum analyzer. (B) Experimental realization of a forced detuning measurement using a low power probe (Pprobe = 0.6 µW) in the strong coupling regime ($g_{{\rm om}}$ = 64 ± 2 MHz, $P_{{\rm pump}}$ = 1.35 ± 0.1 W, $T\leavevmode\nobreak\ \approx$ 4 K). Yellow dashed line represents the optoacoustic eigenvalues. (C) Forced detuning measurement in the weak coupling regime ($g_{{\rm om}}$ = 12 ± 2 MHz, $P_{{\rm pump}}$ = 0.05 ± 0.1 W, $T\leavevmode\nobreak\ \approx$ 4 K). (D) Theoretical simulation of an anti-crossing measurement in a cavity-less optoacoustic system, showing good agreement with the experiment. (E) Same as (D), for the weak coupling case. (F) Spectrum in the strong coupling regime with a red detuned off-resonance probe, (G) with a resonant probe, (H) with a blue detuned off-resonance probe.
@@
FIG. 4: (A) Maxima of the measured strongly coupled optoacoustic modes as a function of pump power. Dark squares: $g_{{\rm om}}$ = 48 ± 2 MHz, medium circles: $g_{{\rm om}}$ = 61 ± 2 MHz, light stars: $g_{{\rm om}}$ = 75 ± 2 MHz. (B) Maxima of the measured strongly coupled optoacoustic modes as a function of pump power according to the theory. Colors represent same input powers as in (A). Dashed line indicates $\Omega_{\rm B}$.
@@ Acknowledgements.
We thank G. Leuchs, M. Sondermann, F. Marquardt, and G. Jara-Schulz for useful discussions and insightful comments. We thank A. Geilen, Z. O. Saffer, L. Fischer, P. Wiedemann, G. Slinkov, J. H. Marines Cabello, R. Jadhav and X. Zeng for their help at different stages of the experiment. L.B.M., C.Z. and B.S. discussed the general ideas and conceived the experiment. L.B.M. built the experimental setups, conceptualized the forced detuning measurement scheme, performed measurements and analyzed the data. C.Z. derived the theoretical models and performed simulations of the experiments. B.S. supervised the project and conceived the project. All authors discussed the results. L.B.M., C.Z. and B.S. wrote the manuscript together.
@@
FIG. 5: (A) The spectra of anti-Stokes wave with pump power of $\rm 0.83\leavevmode\nobreak\ W$ for various wave number $k$, where dotted (dash-dot) lines denote the spectrum of $k$-th anti-Stokes mode with positive (negative) wave number $k$ and the black solid line correspond to the phase-matching case ($k=0$). The dark blue dashed line shows the envelope of spectra for anti-Stokes modes with continuous wave number $k$. The inset indicates that experiment results (grey solid line) agree well with theoretical results (dark blue dashed line). (B) corresponds to the spectra of anti-Stokes wave with pump power of $\rm 1.21\leavevmode\nobreak\ W$.
@@
Fiber type Temperature Length $\Omega_{\rm B}$ $\Gamma_{\rm m}$ $G_{\rm B}$ $g_{\rm B}$
(K) (cm) (GHz) (MHz) (W-1m-1) (W-1m)
@@
FIG. 6: Diagram of the experimental setup used to measure thermal-noise-initiated Brillouin-Mandelstam scattering via heterodyne detection. CW: continuous wave; AOM: acousto-optical modulator; EDFA: erbium-doped fiber amplifier; BPF: band-pass filter, ESA: electrical spectrum analyzer. The three black circles represent a fiber polarization controller.
@@
The results for the experimental study of Brillouin-Mandelstam scattering in the sample at cryogenic temperatures are shown in Fig. 7 and Fig. 8 . As explained in the previous section, not the whole HNLF is at low temperature, as sections of it are outside of the base plate in order to allow splicing for optical access. This results in three regions of the fiber: cold, room temperature and transition between the two. This is reflected in the spectra shown in Fig. 7 (A). Two main peaks are present, the cold one centered at $\Omega_{\rm B}$ = 9.5997 ± 0.0001 GHz and the room temperature one centered at $\Omega_{\rm B}$ = 9.7278 ± 0.0004 GHz. As the pump power is increased, the cold Stokes resonance will enter the stimulated Brillouin-Mandelstam scattering regime (SBS), increasing exponentially and narrowing from 9.91 ± 0.07 MHz to 5.14 ± 0.04 MHz. On the contrary, the anti-Stokes resonance broadens until it splits in two, meaning the strong coupling regime has been reached (Fig. 8 (A)). Having a closer look at the cold part of the spectrum for low input power (Fig. 7 (B) and Fig. 8 (B)), two peaks of different linewidths centered around 9.5997 GHz are present. A narrow 9.91 MHz peak, corresponding to the cryogenic part of the sample, and a broader 77.2 ± 0.7 MHz peak. The source of this signal is the transition regions of the fiber, particularly of temperature around 190 K (Fig. 9 (A)). Around that temperature, the frequency shift is the same as at 4 K (Fig. 9 (B)), but the linewidth is higher (Fig. 9 (C)) [76].
@@
FIG. 7: (A) Normalized Stokes Brillouin response of the sample as a function of pump power. As the pump power is increased, the resonance coming from the 4 K section (peak at 9.5997 GHz) enters the stimulated regime and increases exponentially. The response of the room temperature part (peak at 9.7278 GHz) is only visible for low powers, comparatively. (B) Spectrum of the full Stokes response of the sample at low input power. The signal centered at 9.6 GHz is composed of two peaks, a narrow 9.91 MHz one coming from the 4 K section of the fiber and a broad 77.2 MHz one coming from the transition to room temperature, at 190 K. (C) Spectrum of the full Stokes response of the sample at higher input power. The spectrum is now dominated by the narrow peak coming from the 4 K section of the fiber.
@@
FIG. 8: (A) Normalized Anti-Stokes Brillouin response of the cryogenic sample as a function of pump power. As the pump power is increased, the resonance splits, corresponding to the transition into the strong coupling regime. (B) Spectrum of the full anti-Stokes response of the sample at low input power. Same behavior as described in 7 (B). (C) Spectrum of the full anti-Stokes response of the sample at higher input power. As the pump power increases, the signal coming from the 4 K region of the sample dominates the 9.6 GHz spectrum, showing the strong coupling splitting.
@@
FIG. 9: (A) Stokes response of the sample as a function of sample temperature. The resonance at the top, centered at ($\Omega_{B}$ = 9.72 GHz), corresponds to the room temperature part of the fiber, and it does not shift as its temperature remains constant. On the lower part of the plot, the behavior of the sections of the fiber affected by the variation of temperature is shown. The most pronounced peak corresponds to the fiber section on the base plate of the cryostat. A faint third peak (indicated by the white arrow) is also present, coming from parts of the fiber in the transition. (B) Brillouin shift of two most prominent peaks as a function of sample temperature. The room temperature part of the fiber (red stars) is unaffected by the change in temperature. The temperature-varying peak (blue circles) crosses the 9.6 GHz point twice, at T = 191 K and at T = 4 K. (C) FWHM of the two most prominent peaks as a function of sample temperature. The room temperature part of the fiber (red stars) is unaffected by the change in temperature.
@@
FIG. 10: Diagram of the setup used to measure Brillouin-Mandelstam scattering via a pump-probe experiment with lock-in amplifier detection. EOM: electro-optical modulator; Variable att.: variable attenuator; RF: radio frequency. The remaining abbreviations same as Fig. 6.
@@
FIG. 11: (A) Forced detuning measurement at $T$ = 4 K, $P_{{\rm pump}}$ = 1.35 ± 0.1 W ($g_{{\rm om}}$ = 70 ± 2 MHz). As the system is in the strong coupling regime, it shows a clear avoided crossing. (B) Forced detuning measurement for $T$ = 10 K, same pump power as (A). The avoided crossing is not clearly visible, as the system enters the weak coupling regime. (C) Same as (B) with $T$ = 19 K.

### v2

@@ I Introduction
The particular case of light interacting with traveling acoustic waves, described by Brillouin-Mandelstam scattering, has received interest in recent years, with applications such as sensing [34, 35, 36], lasing [37, 38, 39] or integrated photonics [40]. A wide range of optoacoustic platforms is available, both in material [41, 42, 43, 44, 45] and structural [46, 47, 48, 49, 50, 51] diversity. Waveguides, with their defined direction of propagation, allow to have temporal control over the propagation of optical signals through them, which has in particular been exploited for signal processing [52, 53, 54, 55, 56, 57] and storage [58, 59, 60]. The natural symmetry-breaking of Stokes and anti-Stokes processes has been exploited for active cooling [61, 62], without the need of working in the resolved sideband regime like in resonators [63, 64]. The shape of the spectrum of acoustic modes present in a waveguide is moreover crucially different from that of a resonator. In an optomechanical cavity, narrow discrete resonances are addressed, contrary to the close to continuous spectrum of modes present in a waveguide.
@@
FIG. 1: Cavity-less strong coupling and how to measure it. (a) Sketch of an anti-Stokes Brillouin-Mandelstam scattering process in the strong coupling regime. (b) Dispersion diagram for the different modes involved in an anti-Stokes Brillouin-Mandelstam scattering process. Adapted from [71, 67]. (c) Schematic of a pump-probe experiment. (d) Schematic of a forced detuning experiment.
@@
FIG. 2: Strongly coupled Brillouin response of a $T$ = 4 K fiber. (a) Anti-Stokes Brillouin resonance as function of pump power. (b) Simulation of cavity-less optoacoustic strong coupling [65], showing good agreement with the experiment. (c) and (d) Optical anti-Stokes resonance in the strong coupling regime. (e) Optical anti-Stokes resonance in the weak coupling regime. Experimental data shown as a dark continuous line, theoretical data shown as a light dashed line.
@@
For low input powers, the system is weakly coupled ($g_{{\rm om}}$ = 13 ± 2 MHz) and the anti-Stokes resonance has the shape of a single Lorenzian centered at $\Omega_{\rm B}/2\pi$ = 9.5997 ± 0.0001 GHz (Fig. 2 (e)). As the pump power increases ($g_{{\rm om}}$ = 57 ± 2 MHz), the system enters the strong coupling regime and the resonance splits (Fig. 2 (d)). Continuing to increase the pump power pushes the system deeper into the strong coupling regime ($g_{{\rm om}}$ = 70 ± 2 MHz, Fig. 2 (c)), resulting in a maximum measured split of $\delta_{{\rm split}}$ = 26.7 MHz for $g_{{\rm om}}$ = 71 ± 2 MHz. The shape of these two peaks is asymmetric, with the outer slopes of each peak following the shape of a Lorentzian slope, while the inner slopes are sharper. In order to validate our results, we compare with our theoretical model, which shows good agreement with the experiment (Fig. 2 (b)).
FIG. 3: Forced detuning experiment and results. (a) Diagram of the setup to perform a forced detuning measurement. CW: continuous wave; EDFA: Erbium-doped fiber amplifier; BPF: band-pass filter; E(A)OM: electro(acusto)optical modulator; ESA: electrical spectrum analyzer. (b) Experimental realization of a forced detuning measurement using a low power probe (Pprobe = 0.6 µW) in the strong coupling regime ($g_{{\rm om}}$ = 64 ± 2 MHz, $P_{{\rm pump}}$ = 1.35 ± 0.1 W, $T\leavevmode\nobreak\ \approx$ 4 K). Yellow dashed line represents the optoacoustic eigenvalues. (c) Forced detuning measurement in the weak coupling regime ($g_{{\rm om}}$ = 12 ± 2 MHz, $P_{{\rm pump}}$ = 0.05 ± 0.1 W, $T\leavevmode\nobreak\ \approx$ 4 K). (d) Theoretical simulation of an anti-crossing measurement in a cavity-less optoacoustic system, showing good agreement with the experiment. (e) Same as (d), for the weak coupling case. (f) Spectrum in the strong coupling regime with a red detuned off-resonance probe, (g) with a resonant probe, (h) with a blue detuned off-resonance probe.
@@
FIG. 4: Hybrid optoacoustic mode separation versus input power. (a) Maxima of the measured strongly coupled optoacoustic modes as a function of pump power. Dark squares: $g_{{\rm om}}$ = 48 ± 2 MHz, medium circles: $g_{{\rm om}}$ = 61 ± 2 MHz, light stars: $g_{{\rm om}}$ = 75 ± 2 MHz. (b) Maxima of the measured strongly coupled optoacoustic modes as a function of pump power according to the theory. Colors represent same input powers as in (a). Dashed line indicates $\Omega_{\rm B}$.
@@ Acknowledgements.
We thank G. Leuchs, M. Sondermann, F. Marquardt, and G. Jara-Schulz for useful discussions and insightful comments. We thank A. Geilen, Z. O. Saffer, L. Fischer, P. Wiedemann, G. Slinkov, J. H. Marines Cabello, R. Jadhav and X. Zeng for their help at different stages of the experiment. L.B.M., C.Z. and B.S. discussed the general ideas and conceived the experiment. L.B.M. built the experimental setups, conceptualized the forced detuning measurement scheme, performed measurements and analyzed the data. C.Z. derived the theoretical models and performed simulations of the experiments. B.S. supervised the project and conceived the project. All authors discussed the results. L.B.M., C.Z. and B.S. wrote the manuscript together. The authors acknowledge funding from the Max Planck Research Groups Scheme from the Max Planck Society and the Max Planck School of Photonics. Correspondence and requests for materials should be addressed to Birgit Stiller. The datasets generated during and/or analysed during the current study are available from the corresponding author on reasonable request.
@@
FIG. 5: Simulation results of spectra for the anti-Stokes wave. (a) The spectra of anti-Stokes wave with pump power of $\rm 0.83\leavevmode\nobreak\ W$ for various wave number $k$, where dotted (dash-dot) lines denote the spectrum of $k$-th anti-Stokes mode with positive (negative) wave number $k$ and the black solid line correspond to the phase-matching case ($k=0$). The dark blue dashed line shows the envelope of spectra for anti-Stokes modes with continuous wave number $k$. The inset indicates that experiment results (grey solid line) agree well with theoretical results (dark blue dashed line). (b) corresponds to the spectra of anti-Stokes wave with pump power of $\rm 1.21\leavevmode\nobreak\ W$.
@@
Fiber type Temperature (K) Length (cm) $\Omega_{\rm B}/2\pi$ (GHz) $\Gamma_{\rm m}$ (MHz) $G_{\rm B}$ (W-1m-1) $g_{\rm B}$ (W-1m)
@@
FIG. 6: Thermal-noise-initiated experiment. Diagram of the experimental setup used to measure thermal-noise-initiated Brillouin-Mandelstam scattering via heterodyne detection. CW: continuous wave; AOM: acousto-optical modulator; EDFA: erbium-doped fiber amplifier; BPF: band-pass filter, ESA: electrical spectrum analyzer. The three black circles represent a fiber polarization controller.
@@
The results for the experimental study of Brillouin-Mandelstam scattering in the sample at cryogenic temperatures are shown in Fig. 7 and Fig. 8 . As explained in the previous section, not the whole HNLF is at low temperature, as sections of it are outside of the base plate in order to allow splicing for optical access. This results in three regions of the fiber: cold, room temperature and transition between the two. This is reflected in the spectra shown in Fig. 7 (a). Two main peaks are present, the cold one centered at $\Omega_{\rm B}/2\pi$ = 9.5997 ± 0.0001 GHz and the room temperature one centered at $\Omega_{\rm B}/2\pi$ = 9.7278 ± 0.0004 GHz. As the pump power is increased, the cold Stokes resonance will enter the stimulated Brillouin-Mandelstam scattering regime (SBS), increasing exponentially and narrowing from 9.91 ± 0.07 MHz to 5.14 ± 0.04 MHz. On the contrary, the anti-Stokes resonance broadens until it splits in two, meaning the strong coupling regime has been reached (Fig. 8 (a)). Having a closer look at the cold part of the spectrum for low input power (Fig. 7 (b) and Fig. 8 (b)), two peaks of different linewidths centered around 9.5997 GHz are present. A narrow 9.91 MHz peak, corresponding to the cryogenic part of the sample, and a broader 77.2 ± 0.7 MHz peak. The source of this signal is the transition regions of the fiber, particularly of temperature around 190 K (Fig. 9 (a)). Around that temperature, the frequency shift is the same as at 4 K (Fig. 9 (b)), but the linewidth is higher (Fig. 9 (c)) [76].
@@
FIG. 7: Thermal-noise-initiated measurement of Stokes resonances. (a) Normalized Stokes Brillouin response of the sample as a function of pump power. As the pump power is increased, the resonance coming from the 4 K section (peak at 9.5997 GHz) enters the stimulated regime and increases exponentially. The response of the room temperature part (peak at 9.7278 GHz) is only visible for low powers, comparatively. (b) Spectrum of the full Stokes response of the sample at low input power. The signal centered at 9.6 GHz is composed of two peaks, a narrow 9.91 MHz one coming from the 4 K section of the fiber and a broad 77.2 MHz one coming from the transition to room temperature, at 190 K. (c) Spectrum of the full Stokes response of the sample at higher input power. The spectrum is now dominated by the narrow peak coming from the 4 K section of the fiber.
@@
FIG. 8: Thermal-noise-initiated measurement of anti-Stokes resonances. (a) Normalized Anti-Stokes Brillouin response of the cryogenic sample as a function of pump power. As the pump power is increased, the resonance splits, corresponding to the transition into the strong coupling regime. (b) Spectrum of the full anti-Stokes response of the sample at low input power. Same behavior as described in 7 (b). (c) Spectrum of the full anti-Stokes response of the sample at higher input power. As the pump power increases, the signal coming from the 4 K region of the sample dominates the 9.6 GHz spectrum, showing the strong coupling splitting.
@@
FIG. 9: Stokes Brillouin response versus sample temperature. (a) Stokes response of the sample as a function of sample temperature. The resonance at the top, centered at ($\Omega_{B}/2\pi$ = 9.72 GHz), corresponds to the room temperature part of the fiber, and it does not shift as its temperature remains constant. On the lower part of the plot, the behavior of the sections of the fiber affected by the variation of temperature is shown. The most pronounced peak corresponds to the fiber section on the base plate of the cryostat. A faint third peak (indicated by the white arrow) is also present, coming from parts of the fiber in the transition. (b) Brillouin shift of two most prominent peaks as a function of sample temperature. The room temperature part of the fiber (red stars) is unaffected by the change in temperature. The temperature-varying peak (blue circles) crosses the 9.6 GHz point twice, at T = 191 K and at T = 4 K. (c) FWHM of the two most prominent peaks as a function of sample temperature. The room temperature part of the fiber (red stars) is unaffected by the change in temperature.
@@
FIG. 10: Pump-probe experiment. Diagram of the setup used to measure Brillouin-Mandelstam scattering via a pump-probe experiment with lock-in amplifier detection. EOM: electro-optical modulator; Variable att.: variable attenuator; RF: radio frequency. The remaining abbreviations same as Fig. 6.
@@
FIG. 11: Temperature dependence of strong coupling regime. (a) Forced detuning measurement at $T$ = 4 K, $P_{{\rm pump}}$ = 1.35 ± 0.1 W ($g_{{\rm om}}$ = 70 ± 2 MHz). As the system is in the strong coupling regime, it shows a clear avoided crossing. (b) Forced detuning measurement for $T$ = 10 K, same pump power as (a). The avoided crossing is not clearly visible, as the system enters the weak coupling regime. (c) Same as (b) with $T$ = 19 K.